Atoms of CO 2 gas decay from upper to lower level with a spread in frequency is about
10kHz. Calculate the line shape if the broadening mechanism is found Doppler
broadening ?
1
Expert's answer
2020-05-28T13:42:36-0400

The distribution of the frequencies of the absorbed radiation for Gaussian line shape is subjected to the following law:


"L=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(f-f_0)^2}{2\sigma^2}},"

where


"\sigma=\frac{f_0}{c}\sqrt{\frac{2kT\text{ ln}2}{m}}."

Therefore, the provided information is not enough to calculate the line shape.